[3 pts] A croissant shop has plain croissants, cherry croissants, chocolate croissants, almond croissants, apple croissants, and broccoli croissants. How many ways are there to choose 10 croissants?
Answer #1

here let x1,x2,x3,x4,x5 and x6 are number of plain croissants, cherry croissants, chocolate croissants, almond croissants, apple croissants, and broccoli croissants

therefore x1+x2+x3+x4+x5+x6 =10 where each xi >=0

number of solution of above equation =number of ways to choose 10 croissants

=(10+6-1C6-1) =(15C5) =3003